# Generative AI Enables EEG Super-Resolution via Spatio-Temporal Adaptive Diffusion Learning

### v4

@@ IV-D Classification Evaluation on the Synthetic SR EEG
Fig. 6 shows the comparison results of four classification metrics between LR EEG and synthetic SR EEG. The SR EEG reconstructed by STAD outperforms LR EEG in all metrics, showing that STAD enhance classification performance of EEG-Net. As the channel count of LR EEG increases, the classification performance improvement by SR EEG also increases, indicating better detail reconstruction. This demonstrates that STAD effectively enhance LR EEG spatial resolution, improving downstream classification tasks. Moreover, this provides a novel approach for other EEG signal enhancement and EEG-based clinical diagnosis research.
We also analyzed classification performance from the frequency domain perspective, calculating the PSD of LR and SR EEG in five frequency bands. Each sample corresponds to $c\times 5$ features (where c denotes the number of channels), and classification is performed based on these features. Table II shows classification results for LR and SR EEG across different frequency bands and scaling factors. SR EEG outperforms LR EEG in all frequency bands. As the scaling factor increases, classification accuracy decreases for both but remains higher for SR EEG. Low-frequency bands ($\delta$, $\theta$, $\alpha$) have worse performance than high-frequency bands ($\beta$, $\gamma$), but the average accuracy improvement in low-frequency bands (5.06%) surpasses that in high-frequency bands (4.95%). This suggests that the proposed STAD more Adequately preserves low-frequency information and captures high-frequency variations, improving classifier performance in identifying abnormal EEG features.
@@ IV-E Source Localization Evaluation on the Synthetic EEG
Source localization is a crucial task in neuroscience and clinical applications, as it reveals the active source locations within the brain [51]. To determine the impact of the proposed STAD in enhancing EEG spatial resolution and its effectiveness in improving downstream task performance, we assess it using the source localization task, incorporating both quantitative and qualitative analyses. We utilize localization error [8] for quantitative analysis, defining it as the Euclidean distance between the source localization results and the ground-truth positions. This metric is employed to compare the performance of LR EEG, SR EEG, and HR EEG. For qualitative analysis, we visualize the source localization results on brain maps to provide an intuitive display of the localization performance across different data sources. To ensure fairness, we consistently use eLORETA [52], an unconstrained linear inverse solution that infers the distribution of brain current sources from scalp-recorded EEG signals through precise regularization strategies [53]. By examining both quantitative results and visualization outcomes, we can evaluate whether the proposed method effectively enhances the spatial resolution of EEG.
@@ V Discussion and Conclusion
Experimental results show that while the proposed STAD outperform existing methods in reconstruction performance, larger scaling factors reveal a difference between the synthetic SR EEG and the ground truth, as illustrated in Figure 4. We identified several possible reasons for this: i) With larger scaling factors, the available spatial information for the model is constrained, making it challenging to accurately capture the spatial relationships between channels. ii) The inherently low SNR of EEG increases the difficulty of reconstruction for the model. iii) Significant individual differences between subjects make it challenging for the model to uniformly learn the mapping between HR EEG and LR EEG across different subjects. To address the issue of individual differences, personalized strategies can be designed, which is one of our future research priorities. Despite some limitations, the proposed method has demonstrated its effectiveness in downstream tasks and offers a new approach for enhancing the spatial resolution of EEG signals. In the future, we will further expand the application range of this framework to enhance other brain function signals. Through improving algorithms and models, we plan to apply this technology to more precise clinical diagnosis and treatment scenarios, such as BCI technology and neurofeedback, thus promoting research in neuroscience, cognitive science and other related fields and enhancing the understanding and diagnostic capabilities of neuropsychiatric diseases.

### v5

@@ IV-D Classification Evaluation on the Synthetic SR EEG
Fig. 6 shows the comparison results of four classification metrics between LR EEG and synthetic SR EEG. The SR EEG reconstructed by STAD outperforms LR EEG in all metrics, showing that STAD enhance classification performance of EEG-Net. As the channel count of LR EEG increases, the classification performance improvement by SR EEG also increases, indicating better detail reconstruction. This demonstrates that STAD can effectively enhance spatial resolution of LR EEG, improving downstream classification tasks. Furthermore, this provides a novel approach for other EEG signal enhancement and EEG-based clinical diagnosis research.
@@
We also analyzed classification performance from the frequency domain perspective, calculating the PSD of LR and SR EEG in five frequency bands. Each sample corresponds to $c\times 5$ features (where c denotes the number of channels), and classification is performed based on these features. Table II shows classification results for LR and SR EEG across different frequency bands and scaling factors. SR EEG outperforms LR EEG in all frequency bands. As the scaling factor increases, classification accuracy decreases for both but remains higher for SR EEG. Low-frequency bands ($\delta$, $\theta$, $\alpha$) have worse performance than high-frequency bands ($\beta$, $\gamma$), but the average accuracy improvement in low-frequency bands (5.06%) surpasses that in high-frequency bands (4.95%). This finding can be attributed to the fact that low-frequency bands provide more accessible yet less comprehensive information, while high-frequency bands, despite being richer in detail, are more challenging to harness effectively due to their complexity and susceptibility to noise. These results suggest that the proposed STAD framework is adept at preserving low-frequency information while also capturing high-frequency variations, thereby enhancing the classifier’s ability to identify abnormal EEG features with greater accuracy.
@@ IV-E Source Localization Evaluation on the Synthetic EEG
Source localization is a crucial task in neuroscience and clinical applications, as it reveals the active source locations within the brain [54]. To determine the impact of the proposed STAD in enhancing EEG spatial resolution and its effectiveness in improving downstream task performance, we assess it using the source localization task, incorporating both quantitative and qualitative analyzes. Specifically, we first use the subject’s MRI data to construct an individualized forward model. Then, we employ the eLORETA [55] as inverse solver algorithm which infers the distribution of brain current sources from scalp-recorded EEG signals through precise regularization strategies [56]. Furthermore, we use localization error [9] as the evaluation metric, defining it as the Euclidean distance between the source localization results and the ground-truth positions. This metric is employed to compare the performance of LR EEG, SR EEG, and HR EEG. For qualitative analysis, we visualize the source localization results on brain maps to provide an intuitive display of the localization performance across different data sources.
@@ V Discussion and Conclusion
Experimental results show that while the proposed STAD outperform existing methods in reconstruction performance, larger scaling factors reveal a difference between the synthetic SR EEG and the ground truth, as illustrated in Fig. 4. We identified several possible reasons for this: i) With larger scaling factors, the available spatial information for the model is constrained, making it challenging to accurately capture the spatial relationships between channels. ii) The inherently low SNR of EEG increases the difficulty of reconstruction for the model. iii) Significant individual differences between subjects make it challenging for the model to uniformly learn the mapping between HR EEG and LR EEG across different subjects. To address the issue of individual differences, personalized strategies can be designed, which is one of our future research priorities. Despite some limitations, the proposed method has demonstrated its effectiveness in downstream tasks and offers a new approach for enhancing the spatial resolution of EEG signals.
Figure 9: The framework for Constructing correlation mechanisms among different brain functional modalities using the generative AI paradigm.
In our future work, we aim to broaden the applicability of generative AI to enhance other brain function signals. As shown in Fig. 9, we plan to employ this data-driven approach to address a broad range of neural signal tasks, including deep spatial super-resolution reconstruction of invasive EEG (iEEG), brain semantic decoding and the establishment of correlation mechanism among neural electrical signals, neural magnetic signals, and brain blood oxygenation signals, like fMRI and functional Near-Infrared Spectroscopy (fNIRS). Moreover, we will actively advance the integration of generative AI into real-world scenarios such as brain-computer interfaces, neurofeedback, and telemedicine. We anticipate that this diversified integration will not only propel research endeavors in neuroscience, cognitive science, and related fields but also enhance our understanding and diagnostic accuracy of neuropsychiatric disorders.
VI Acknowledgment
This work was supported by the National Key Research and Development Program of China under Grant 2023YFC2506902, National Natural Science Foundations of China under Grant 62172403, 12326614 and U2241210, and Strategic Priority Research Program of Chinese Academy of Sciences under Grant XDB38040200.